U= (-1,-7)
W= (3,1)
2U+(-3)w=
1
Expert's answer
2020-04-22T17:44:59-0400

"2U+(-3)W=2(-1,-7)+(-3)(3,1)="

"=(2\cdot(-1)+(-3)\cdot 3,2\cdot(-7)+(-3)\cdot 1)="

"=(-11,-17)"

Answer: "(-11,-17)"